OpenStudy (anonymous):

Given Euler's formula, prove that cos(x)=(e(^x)+e(^-x))/2 and sin(x)=(e(^x)-e(^-x))/2i

OpenStudy (anonymous):

\[\cos(x)=(e^x+e^{-x})/2\] and \[\sin(x)=(e^x-e^{-x})/2i\] more neatly

OpenStudy (earthcitizen):

\[e ^{j \theta} =\cos( \theta)+jsin(\theta)\]

OpenStudy (earthcitizen):

\[\cos(x)=(e ^{x}/2)+(e ^{-x}/2)\]

OpenStudy (anonymous):

Ah, yes! It's easier when you express e^-x, e^x and cos/sin(x) as an infinite series. I think I get it now.

OpenStudy (y2o2):

I think this can't be true bec. \[iSin(x) = {(e^x - e^{-x} ) \over 2}\] \[\cos(x) + isin(x) = {{e^x + e^{-x}}\over 2} + {{e^x - e^{-x}}\over 2} = {2e^x \over 2} = e^x \] which is wrong bec cos(x) + isin(x) = e^(ix) Not e^x

OpenStudy (anonymous):

Bah! Typo- sorry for utterly wasting your time!

OpenStudy (earthcitizen):

\[e ^{jx}\] not e^x